# Dynamics of a compressible gas injected into a confined porous layer

### v2

@@ 1 Introduction
Underground gas storage plays a crucial role in addressing anthropogenic climate change, both by sequestering greenhouse gas emissions prior to atmospheric release and by enabling the transition to cleaner energy sources. The first approach is well-established through carbon capture and storage, where large quantities of CO2 are injected underground for long-term containment. More recently, underground hydrogen storage has gained significant attention for its potential to enable large-scale integration of renewable energy into the power grid (heinemann_enabling_2021; zivar_underground_2021; muhammed_review_2022; lord_geologic_2014; tarkowski_underground_2019). The concept is to use surplus renewable energy generated during the summer months to produce hydrogen via electrolysis, store it in underground formations, and subsequently recover it during the winter when demand exceeds supply. Hydrogen is a particularly attractive candidate for this large-scale energy storage owing to its high energy density and the fact that its combustion produces minimal harmful emissions (zivar_underground_2021). Proposed underground storage sites for gas include depleted reservoirs, aquifers, and salt caverns, all of which are typically saturated with brine. Unlike CO2 storage, where the gas is intended for permanent sequestration, hydrogen storage requires recoverability, making accurate predictions of gas plume evolution crucial. Due to its exceptionally low viscosity relative to brine, injected hydrogen may spread rapidly along the top of the brine within an aquifer, forming a thin gas layer that extends far from the injection site (hagemann_mathematical_2015). This behaviour not only reduces storage efficiency but also increases the risk of hydrogen loss.
@@
huppert_gravity-driven_1995 examined the exchange of two confined incompressible fluids with differing densities and pressures, but equal viscosities, between two aquifers. By deriving a similarity solution, they showed that buoyancy transports the denser fluid along the lower boundary and the lighter fluid along the upper boundary, while an imposed pressure gradient drives a net flow. Subsequent research expanded on this framework to incorporate fluids with different viscosities, highlighting how the viscosity contrast influences spreading behaviour (nordbotten_similarity_2006; pegler_fluid_2014; zheng_flow_2015). nordbotten_similarity_2006 generalized Huppert & Woods’ (huppert_gravity-driven_1995) study to investigate the effect of viscosity contrasts in an axisymmetric geometry. pegler_fluid_2014 conducted a comprehensive study of flow behaviour in a two-dimensional channel. Their findings revealed that, at early times, spreading is primarily buoyancy-driven and can be approximated by the classical porous-medium equation of barenblatt_unsteady_1952. At later times, low-viscosity injected fluid forms a thin layer along the upper boundary, with buoyancy becoming less dominant compared to viscous effects. By deriving a large-time similarity solution, they demonstrated that the movement of the upper contact line is controlled by the strength of the source. In this case, since the flow is incompressible, the length of the channel determines the pressure needed to drive the motion but does not affect the spreading dynamics. These theoretical predictions were validated through laboratory experiments in which freshwater was injected into a saltwater-saturated porous medium, confirming that the similarity solution accurately describes the interface evolution.
@@
When the injected gas is compressible, however, the dynamics change: the source can in principle increase the gas pressure without advancing the gas-liquid interface, and the subsequent propagation of the upper contact line may be influenced by the length of the channel. In long channels, the viscous pressure drop between source and outlet may lead to spatial variations in pressure, which could in turn lead to noticeable variations in gas density along the channel. These variations could affect the mass flow rate and, consequently, influence the rate of gas spreading. Despite the potential for compressibility to significantly alter the flow dynamics, its effects in this context remain under-explored. While some studies have addressed compressibility, most have focused on fluid-structure interactions, particularly the impact of pressure buildup on the rock matrix integrity (mathias_approximate_2009) and fluid migration in neighboring reservoirs (jenkins_impact_2019). A common approach to coupling the mechanics of the pore structure with pressure build-up during injection is to introduce rock compressibility, defined as $c_{r}=({1}/{\phi}){\mathrm{d}\phi}/{\mathrm{d}p^{*}}$, where $\phi$ is the porosity and $p^{*}$ is the pressure. This parameter quantifies how the pore space deforms in response to pressure changes and is valid under conditions of constant vertical stress and negligible lateral strain (jenkins_impact_2019).
@@
In this paper, we extend the framework of pegler_fluid_2014 for injection of a gas into a confined porous channel containing brine (figure 1) to investigate how gas compressibility interacts with buoyancy and viscous effects in both fluids to determine the spreading dynamics. Using a long-wave approximation, we derive two coupled nonlinear evolution equations governing the gas/liquid interface height and the gas pressure field. These are mass conservation equations, with the gas pressure serving as a proxy for gas density via the equation of state. The viscous pressure drop ahead of the upper contact line is captured by a boundary condition, while two kinematic conditions determine the motion of the contact lines. The model admits several distinguished asymptotic limits that reveal the roles of different mechanisms across parameter space; we derive a set of reduced models that capture dominant balances. In particular, we estimate key quantities of practical interest, including the pressure scale, pressure rise time, and the breakthrough time. In §2, we present the governing equations and describe the scaling argument used to derive the long-wave model. The system is characterized by three primary dimensionless parameters, and we simplify the model in various sub-limits, identifying relevant scaling relationships. In §LABEL:sec:results, we present numerical solutions of the full model, illustrating how the key features of compressible flow emerge and comparing them with predictions from asymptotic analysis. Finally, in §LABEL:sec:discussion, we revisit the dominant balances in each regime, interpret the underlying physical mechanisms, and discuss the broader implications for hydrogen storage and related gas-injection technologies.
@@ 2.1 Governing equations
We consider a two-dimensional model of fluid displacement driven by the injection of a compressible gas into a planar horizontally-confined porous medium with depth $H$ and length $L$, as illustrated in figure 1. Initially, the medium is predominantly saturated with a liquid, with only a small amount of gas present. The length of the region initially occupied by gas is denoted by $L_{g0}$; the aspect ratio of this region is assumed small, i.e. $H/L_{g0}\ll 1$. (We do not seek here to model the initial formation of the gas bubble.) The permeability of the medium is taken to be uniform, with value $k_{0}$.
At time $t^{*}=0$, gas is injected into the channel from a line source located at $\boldsymbol{x}^{*}~\equiv(x^{*},y^{*})=\boldsymbol{x}^{*}_{0}$. Variations in the injection rate are governed by the function $q\,\mathcal{Q}(\omega{t^{*}})$, where $1/\omega$ represents the timescale of injection, $q$ denotes the source strength per unit width of the channel, and $\mathcal{Q}$ is a dimensionless function. The pressure at the outlet ($x^{*}=L$) is taken to be hydrostatic with a baseline value of $p_{g0}$. Using a sharp-interface model, disregarding capillary or miscibility effects, the fluids are segregated into two distinct phases and the domain is divided into four regions (figure 1): the gas occupies regions I and IIg, while the liquid occupies regions IIw and III. We use the subscripts $g$ and $w$ to represent quantities in the gas and liquid (water) phase respectively. We revisit all modelling assumptions in §4.
@@
[…]
The liquid is assumed to be incompressible with constant density ${\rho}_{w}$. The initial gas and liquid volumes are assumed to be in thermal equilibrium with the surrounding pore matrix. Gas compression is treated as isothermal, an assumption that is valid when the thermal inertia of the solid pore matrix is much greater than that of the gas (kushnir_temperature_2012) At a fixed temperature, we define $\rho_{g0}$ to be the equilibrium gas density when at the outlet pressure $p_{g0}$. Then, using $p_{g0}$ and $\rho_{g0}$ as a reference pressure and a reference density, we write the equation of state for the gas (relating density $\rho^{*}_{g}$ to pressure $p^{*}_{g}$) and the equilibrium speed of sound in the gas ($c$) as

### v3

@@ 1 Introduction
Underground gas storage plays a crucial role in addressing anthropogenic climate change, both by sequestering greenhouse gas emissions prior to atmospheric release and by enabling the transition to cleaner energy sources. The first approach is well-established through carbon capture and storage, where large quantities of CO2 are injected underground for long-term containment. More recently, underground hydrogen storage has gained significant attention for its potential to enable large-scale integration of renewable energy into the power grid (heinemann_enabling_2021; zivar_underground_2021; muhammed_review_2022; lord_geologic_2014; tarkowski_underground_2019). The concept is to use surplus renewable energy generated during the summer months to produce hydrogen via electrolysis, store it in underground formations, and subsequently recover it during the winter when demand exceeds supply. Hydrogen is a particularly attractive candidate for this large-scale energy storage owing to its high energy density and the fact that its combustion or use in a fuel cell produces minimal harmful emissions (zivar_underground_2021). Proposed underground storage sites for gas include depleted reservoirs, aquifers, and salt caverns, all of which are typically saturated with brine. Unlike CO2 storage, where the gas is intended for permanent sequestration, hydrogen storage requires recoverability, making accurate predictions of gas plume evolution crucial. Due to its exceptionally low viscosity relative to brine, injected hydrogen may spread rapidly along the top of the brine within an aquifer, forming a thin gas layer that extends far from the injection site (hagemann_mathematical_2015). This behaviour not only reduces storage efficiency but also increases the risk of hydrogen loss.
@@
huppert_gravity-driven_1995 examined the exchange of two confined incompressible fluids with differing densities and pressures, but equal viscosities, between two aquifers. By deriving a similarity solution, they showed that buoyancy transports the denser fluid along the lower boundary and the lighter fluid along the upper boundary, while an imposed pressure gradient drives a net flow. Subsequent research expanded on this framework to incorporate fluids with different viscosities, highlighting how the viscosity contrast influences spreading behaviour (nordbotten_similarity_2006; pegler_fluid_2014; zheng_flow_2015). nordbotten_similarity_2006 generalised Huppert & Woods’ (huppert_gravity-driven_1995) study to investigate the effect of viscosity contrasts in an axisymmetric geometry. pegler_fluid_2014 conducted a comprehensive study of flow behaviour in a two-dimensional channel. Their findings revealed that, at early times, spreading is primarily buoyancy-driven and can be approximated by the classical porous-medium equation of barenblatt_unsteady_1952. At later times, low-viscosity injected fluid forms a thin layer along the upper boundary, with buoyancy becoming less dominant compared to viscous effects. By deriving a large-time similarity solution, they demonstrated that the movement of the upper contact line is controlled by the strength of the source. In this case, since the flow is incompressible, the length of the channel determines the pressure needed to drive the motion but does not affect the spreading dynamics. These theoretical predictions were validated through laboratory experiments in which freshwater was injected into a saltwater-saturated porous medium, confirming that the similarity solution accurately describes the interface evolution.
@@
When the injected gas is compressible, however, the dynamics change: the source can in principle increase the gas pressure without advancing the gas-liquid interface, and the subsequent propagation of the upper contact line may be influenced by the length of the channel. In long channels, the viscous pressure drop between source and outlet may lead to spatial variations in pressure, which could in turn lead to noticeable variations in gas density along the channel. These variations could affect the mass flow rate and, consequently, influence the rate of gas spreading. Despite the potential for compressibility to significantly alter the flow dynamics, its effects in this context remain under-explored. While some studies have addressed compressibility, most have focused on fluid-structure interactions, particularly the impact of pressure buildup on the rock matrix integrity (mathias_approximate_2009) and fluid migration in neighbouring reservoirs (jenkins_impact_2019). A common approach to coupling the mechanics of the pore structure with pressure build-up during injection is to introduce rock compressibility, defined as $c_{r}=({1}/{\phi}){\mathrm{d}\phi}/{\mathrm{d}p^{*}}$, where $\phi$ is the porosity and $p^{*}$ is the pressure. This parameter quantifies how the pore space deforms in response to pressure changes and is valid under conditions of constant vertical stress and negligible lateral strain (jenkins_impact_2019).
@@
In this paper, we extend the framework of pegler_fluid_2014 for injection of a gas into a confined porous channel containing brine (figure 1) to investigate how gas compressibility interacts with buoyancy and viscous effects in both fluids to determine the spreading dynamics. Using a long-wave approximation, we derive two coupled nonlinear evolution equations governing the gas/liquid interface height and the gas pressure field. These are mass conservation equations, with the gas pressure serving as a proxy for gas density via the equation of state. The viscous pressure drop ahead of the upper contact line is captured by a boundary condition, while two kinematic conditions determine the motion of the contact lines. The model admits several distinguished asymptotic limits that reveal the roles of different mechanisms across parameter space; we derive a set of reduced models that capture dominant balances. In particular, we estimate key quantities of practical interest, including the pressure scale, pressure rise time, and the breakthrough time. In § 2, we present the governing equations and describe the scaling argument used to derive the long-wave model. The system is characterised by three primary dimensionless parameters, and we simplify the model in various sub-limits, identifying relevant scaling relationships. In § 3, we present numerical solutions of the full model, illustrating how the key features of compressible flow emerge and comparing them with predictions from asymptotic analysis. Finally, in § 4.1, we revisit the dominant balances in each regime, interpret the underlying physical mechanisms, and discuss the broader implications for hydrogen storage and related gas-injection technologies.
@@ 2.1 Governing equations
We consider a two-dimensional model of fluid displacement driven by the injection of a compressible gas into a planar horizontally-confined porous medium with height $H$ and length $L$, as illustrated in figure 1. Initially, the medium is predominantly saturated with a liquid, with only a small amount of gas present. The length of the region initially occupied by gas is denoted by $L_{g0}$; the aspect ratio of this region is assumed small, i.e. $H/L_{g0}\ll 1$. (We do not seek here to model the initial formation of the gas bubble.) The permeability of the medium is taken to be uniform, with value $k_{0}$.
At time $t^{*}=0$, gas is injected into the channel from a line source located at $\boldsymbol{x}^{*}~\equiv(x^{*},y^{*})=\boldsymbol{x}^{*}_{0}$. Variations in the injection rate are governed by the function $q\,\mathcal{Q}(\omega{t^{*}})$, where $1/\omega$ represents the timescale of injection, $q$ denotes the source strength per unit width of the channel, and $\mathcal{Q}$ is a dimensionless function. The pressure at the outlet ($x^{*}=L$) is taken to be hydrostatic with a baseline value of $p_{g0}$. Using a sharp-interface model, disregarding capillary or miscibility effects, the fluids are segregated into two distinct phases and the domain is divided into four regions (figure 1): the gas occupies regions I and IIg, while the liquid occupies regions IIw and III. We use the subscripts $g$ and $w$ to represent quantities in the gas and liquid (water) phase respectively. We revisit all modelling assumptions in § 4.1.
[…]
@@
The liquid is assumed to be incompressible with constant density ${\rho}_{w}$. The initial gas and liquid volumes are assumed to be in thermal equilibrium with the surrounding pore matrix. Gas compression is treated as isothermal, an assumption that is valid when the thermal inertia of the solid pore matrix is much greater than that of the gas (kushnir_temperature_2012). At a fixed temperature, we define $\rho_{g0}$ to be the equilibrium gas density when at the outlet pressure $p_{g0}$. Then, using $p_{g0}$ and $\rho_{g0}$ as a reference pressure and a reference density, we write the equation of state for the gas (relating density $\rho^{*}_{g}$ to pressure $p^{*}_{g}$) and the equilibrium speed of sound in the gas ($c$) as